OpenStudy (lcnettech):

please help .What is the GCF of 24 and 44?

OpenStudy (lcnettech):

@sammixboo

OpenStudy (lcnettech):

@sammixboo

OpenStudy (anonymous):

Let us say you have worked out the factors of two numbers: Example: Factors of 12 and 30 Factors of 12 are 1, 2, 3, 4, 6 and 12 Factors of 30 are 1, 2, 3, 5, 6, 10, 15 and 30 Then the common factors are those that are found in both lists: Notice that 1, 2, 3 and 6 appear in both lists? So, the common factors of 12 and 30 are: 1, 2, 3 and 6 It is a common factor when it is a factor of two or more numbers. (It is then "common to" those numbers.

OpenStudy (anonymous):

now with that in mind ,make a list of the factors.

OpenStudy (lcnettech):

but what if you can only use one number

OpenStudy (lcnettech):

as the answer

OpenStudy (anonymous):

then in the example I gave you 6 would be the GCF

OpenStudy (anonymous):

The "Greatest Common Factor" is the largest of the common factors

OpenStudy (lcnettech):

so the answer to my question would be 4?

OpenStudy (anonymous):

24 :1,3,4,6,8,12 ,24 44:1 44 2 22 4 10 so yes 4 is the GCF

OpenStudy (anonymous):

*44 : 1,2,4,10,22,44

OpenStudy (anonymous):

A more general way of doing this problem and other problems dealing with big numbers is to write the prime factorization of each number to find the GCF \[ 24 = 2^3\times 3\\ 44 = 2^2 \times 11 \] You can see that \( 2^2=4\) is the GCF

OpenStudy (lcnettech):

can you help with a few more

OpenStudy (anonymous):

I'll try to :)

OpenStudy (lcnettech):

What is the LCM of 15 and 60

OpenStudy (anonymous):

A common multiple is a number that is a multiple of two or more numbers. The common multiples of 3 and 4 are 0, 12, 24, .... The least common multiple (LCM) of two numbers is the smallest number (not zero) that is a multiple of both.

OpenStudy (anonymous):

To practice problems about prime factorizations visit http://www.saab.org/mathdrills/primef.cgi

OpenStudy (anonymous):

15 and 60 would be ....

OpenStudy (anonymous):

Also http://www.saab.org/mathdrills/lcmgcd.cgi

OpenStudy (lcnettech):

3?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

What grade are you in ?

OpenStudy (lcnettech):

6th

OpenStudy (anonymous):

oh alright . This isnt for a test right?

OpenStudy (lcnettech):

no

OpenStudy (anonymous):

okay :) Is that the last question ?

OpenStudy (lcnettech):

yep :)

OpenStudy (lcnettech):

thanks

OpenStudy (anonymous):

No problem

OpenStudy (lcnettech):

how do i find what the least common denominator of a fraction is?

OpenStudy (anonymous):

Look at theses two sites :) 1) http://www.mathsisfun.com/least-common-denominator.html 2) http://www.mathsisfun.com/least-common-multiple-tool.html Hope they help :)